### v2

@@ Abstract
Table structure recognition is a key task in document analysis. However, the geometric deformation in deformed tables causes a weak correlation between content information and structure, resulting in downstream tasks not being able to obtain accurate content information. To obtain fine-grained spatial coordinates of cells, we propose the OG-HFYOLO model, which enhances the edge response by Gradient Orientation-aware Extractor, combines a Heterogeneous Kernel Cross Fusion module and a scale-aware loss function to adapt to multi-scale objective features, and introduces mask-driven non-maximal suppression in the post-processing, which replaces the traditional bounding box suppression mechanism. Furthermore, we also propose a data generator, filling the gap in the dataset for fine-grained deformation table cell spatial coordinate localization, and derive a large-scale dataset named Deformation Wired Table (DWTAL). Experiments show that our proposed model demonstrates excellent segmentation accuracy on all mainstream instance segmentation models. The dataset and the source code are open source: https://github.com/justliulong/OGHFYOLO.
@@ 1 Introduction
In the digital age of increasing information, tables, as the core carrier of structured data, widely carry key information such as financial reports, education and learning, and scientific experimental data. At the same time, with the popularization of scanning, photography and other technologies, the scene of electronic documents of tables is becoming more and more complex. Especially in the case of complex and changing conditions of illumination, angle and scene conditions, the scanned table images or photographed table images may have various deformations. These physical deformations (such as bending, perspective distortion, and folds ) bring great challenges to table structure recognition technology.
The table structure recognition technology [1], which aims at reconstructing the row-column topology and semantic information of a table from an image, has made significant progress through deep learning. Although there are many end-to-end models, which realize the transition from image to structure sequence by directly generating HTML or LaTeX table structure sequences [2, 3], but the black-box feature of end-to-end models makes the intermediate steps uncontrollable. In contrast, the non-end-to-end strategy, a modular approach that decouples cell localization and structural inference, provides a relatively more transparent solution for the table structure recognition task, which also creates the need for cell space coordinates localization. This topic takes the acquisition of cell space coordinates as the research content and studies a method that can effectively extract cell space coordinates in severely deformed tables.
The spatial coordinate localization of table cells can be primarily divided into two methods. One is object detection based on cell contours, which effectively detects most cells with mild deformation, but in the case of serious deformation, this detection method may lose the key information of the cell content, resulting in weak correlation between the cell’s content information and the structure, which brings difficulties to downstream tasks (such as cell content acquisition). The other is text box based segmentation[4, 5], mainly applied in wireless non-deformed tables but has not been widely used in wired deformed tables. So there is still a gap in the acquisition of spatial coordinates of wired table cells with severe deformation. Simply using keypoint-based object detection cannot solve the problem of deformed images (as shown in Figure 1(b)). To address this gap, this study proposes an instance segmentation-based framework (Fig. 1(c)) that achieves finer-grained spatial coordinate localization, which can effectively provide more accurate and rich information for downstream tasks.
@@
Figure 1: (1(a)) shows a deformed table under natural scene, (1(b)) demonstrates the detection results by contour corner-points, and (1(c)) presents the instance segmentation-based contour, as shown, instance segmentation maintains complete content in deformed cells, providing richer information for downstream tasks like content restoration.
@@
On one hand, there are still some gaps in the datasets for deformed wire table recognition tasks. According to public literature, only a few large-scale open-source datasets like WTW[6] and TAL-OCR[7] target deformed table scenarios, but their annotation granularity is difficult to support the training of pixel-level segmentation models. To address this, we design a data generator that can generate new datasets based on existing datasets. Specifically, we select mildly deformed tables from TAL-OCR and WTW, convert their labels into finer-grained segmentation annotations, and feed this annotated data into the data generator to generate a large amount of deformed table data, effectively supporting instance segmentation model training. Based on some data from the open-source datasets TAL-OCR and WTW, as well as 150 images collected and labeled offline, we derive a large-scale dataset through the data generator and divide it into two independent parts according to the sources: Deformed Wire Table for Small (DWTAL-s) with 8,765 simpler tables mainly derived from TAL-OCR, and Deformed Wire Table for Large (DWTAL-l) containing 19,520 complex tables mainly expanded from WTW. Subsequent experiments focus on these two datasets.
On the other hand, two key challenges emerge when using instance segmentation for the spatial localization of cells in deformed wire tables: Firstly, the highly dense arrangement of table cells and their shared contour lines between adjacent cells make boundary information extraction difficult. Secondly, the existence of merged cells leads to great variation in cell scales, including excessively wide-flat or narrow-tall cells, further complicating the spatial localization of cells.
Instance segmentation is dedicated to achieving pixel-level object separation and semantic parsing, and its method reflects the exploration of the balance between accuracy and efficiency in computer vision. Early studies mainly relied on two-stage frameworks. A representative work, Mask R-CNN [8], achieved highly accurate mask predictions in complex scenes by generating region proposal and extracting features.However, such method is difficult to meet real-time application requirements due to the inherent latency of multi-stage computational processes. With the rise of the single-stage method, researchers have discarded the region proposal mechanism, and instead directly generate object locations and masks through intensive prediction of the global feature map. Nevertheless, single-stage models still face significant challenges in terms of object outlines complexity and dense object differentiation, and need to achieve accuracy breakthroughs through innovations in feature expression mechanisms.
@@
In order to solve the challenges of dense objectives and large-scale changes brought by the derived dataset, we firstly propose a Gradient Orientation-aware Extractor (GOE), which integrates the gradient direction and gradient strength of the image texture into the YOLO-structured network as the learning features and can capture the boundary information of each cell more adequately. Meanwhile, the model combines the bottleneck structure [9] and Heterogeneous Kernel Select Protocol [10] (HKSP) in the feature fusion stage, which can more effectively fuse the features of different scales and improve the detection effect of the model in the case of large-scale changes. Additionally, a scale-aware loss function is designed for segmentation tasks, which adaptively sets higher loss weights to smaller objectives to improve model robustness to scale variations. In the post-processing stage, the traditional instance segmentation model usually follows the non-maximal suppression mechanism based on the bounding box; however, due to the densely arranged deformed cells with complex shapes, the bounding box will have a high degree of overlapping phenomenon, which will lead to the boundary box-based NMS incorrectly suppressing the correct objective. For this reason, we introduce the mask-based NMS, which directly performs suppression judgment based on the intersection over union of the predicted masks. The main contributions of this work are summarized as follows:
@@
  • To address objective omission caused by densely arranged cells in deformed tables, we propose Gradient Orientation-aware Extractor (GOE) that enhances boundary discrimination capability for densely arranged cells. The GOE integrates dual cues of gradient intensity and orientation: During feature extraction, directional filters decouple the image gradient field to quantify intensity distribution in edge regions, preventing isotropic diffusion of conventional convolution kernels along edges. Concurrently, an orientation-aware attention mechanism is designed to encode gradient directions as channel components, guiding the deep layers of the network to focus on discriminative regions of deformation boundaries. we also proposes a Heterogeneous Kernel Cross Fusion (HKCF) module for feature fusion stage. This module dynamically adapts feature representation for multi-scale objectives through the Heterogeneous Kernel Select Protocol (HKSP), deploying asymmetric convolution kernels in parallel. By integrating cross-layer feature interaction mechanisms, the HKCF explicitly captures differentiated spatial patterns of horizontally wide cells and vertically narrow cells.
@@
  • we further designs a scale-aware loss function that assigns adaptive weights to objective based on their scales, enhancing the model’s sensitivity to smaller deformed cells while suppressing feature redundancy in larger ones. Furthermore, to prevent valid cells from being erroneously suppressed during post-processing due to complex shapes and dense arrangements of deformed cells, we abandon conventional bounding box-based non-maximum suppression (NMS) and introduce a mask-driven NMS post-processing operation.
@@
  • To address the scarcity of fine-grained annotated datasets for deformed tables, we designed a data generator that expand existing datasets. We first select mildly deformed tables from available datasets, converting their labels into finer-grained segmentation annotations. These curated samples are then processed by the generator to derive two datasets with varying difficulty levels and scales: DWTAL-s and DWTAL-l.
@@ 2.1 Spatial coordinate localization of table cells
The spatial coordinate localization of table cells, as an upstream task in table structure recognition, plays a crucial role in subsequent table structure recognition and parsing tasks. To address cell localization, researchers have proposed various approaches. For instance, Prasad et al. [4] introduced CascadeTabNet, which treats text boxes in tables as objectives and formulates text box detection as an instance segmentation task. Built upon Cascade Mask R-CNN, the model extracts text box masks. This represents the first application of instance segmentation algorithms in table structure recognition. However, their experiments were limited to non-deformable tables, and the impact of empty cells on downstream tasks was not thoroughly discussed.
Following a similar concept, Qiao Liang et al. [5] developed the LGPMA model, which integrates global and local information through a soft feature pyramid while also employing instance segmentation for text box detection. Unlike CascadeTabNet, LGPMA incorporates a dedicated search algorithm for empty cells. Regrettably, this model was not evaluated on deformable tables either. To address spatial coordinate extraction in deformable tables, Cycle-CenterNet [6] proposed a contour corner point-based object detection approach that predicts four contour points by detecting cell centers. While advancing deformable table recognition, this method faces limitations: severe boundary curvature significantly increases regression difficulty for corner points, and critical information in certain cells may be lost when using coarse-grained object detection frameworks.
These challenges highlight the need for finer-grained techniques to preserve essential cell information. Motivated by this requirement, our work introduces instance segmentation technology into spatial coordinate localization for deformable tables. By leveraging the pixel-level precision of instance segmentation, our approach achieves more refined spatial coordinate extraction, thereby better addressing the challenges of cell localization in deformable table structures.
@@ 2.2 Existing dataset
The development of table structure recognition has been propelled by numerous open-source datasets. Early datasets like UNLV [11] and ICDAR-2013 [12] were primarily designed for evaluating traditional methods, containing limited samples (typically under 1,000 images) and lacking spatial coordinate annotations for table cells. These limitations reduce their utility for modern deep learning-based modular table recognition pipelines. As deep learning demands increasingly larger datasets, researchers like Xu Zhong and Minghao Li introduced large-scale datasets such as PubTabNet [2] and TableBank [13]. However, these datasets focus on structural annotations using HTML or LaTeX sequences while neglecting spatial coordinate labeling. Similarly, large-scale datasets like FinTabNet [14] and SciTSR [15] include cell coordinates and row-column relationships but primarily data derived from structured digital documents, such as PDFs or documents exported from LaTeX, which exhibit a high degree of standardization, failing to address deformable table recognition.
Although ICDAR-2019 [16] attempted to bridge this gap by introducing scanned archival documents, its small scale (750 images) and limited deformation types remain insufficient. The CamCap [17] dataset, explicitly designed for rule-based algorithms, contains only 85 deformed tables—far too small to support data-driven deep learning models. Until the WTW[6] dataset was proposed which captures tables in diverse natural scenes with complex backgrounds and extensive deformations. However, WTW employs coarse-grained spatial coordinate annotations based on four contour corner points, which struggle with severe deformations. Similarly, the education-focused TAL-OCR [7] dataset, collected from real-world photography, includes moderate deformations but also lacks fine-grained instance segmentation-level spatial annotations for cells. None of these datasets adequately address the challenges of recognizing heavily deformed table structures.
@@ 3 Data generator and derivative dataset
This section first introduces the implementation of the data generator, and then briefly presents the dataset DWTAL derived from the data generator.
@@ 3.1 Data generator
Deep learning is inherently data-driven, where the scale and quality of datasets significantly influence model performance. To generate a sufficiently large and high-quality dataset, we propose a data generator designed to create distorted images and corresponding labels from existing images and annotations. The distortions are carefully crafted to simulate real-world deformation scenarios. Specifically, the generator incorporates two distortion methods: wave warping and cylindrical warping. Additionally, it adjusts image brightness based on the original illumination to mimic natural lighting variations in real-world photography. The following three subsections detail the implementation of these two distortion techniques and the strategy for brightness adjustment.
@@ 3.1.1 Wave distortion transformation
Wave warping is a wave-like warping of images based on trigonometric functions. This technique is designed to mimic distortions caused by attaching images to flexible curved surfaces (e.g., corrugated pipes or folded fabrics), which induce sine or cosine function-based warping, or to imitate human-induced distortions. Such warping introduces severe deformations that significantly challenge table structure recognition. The specific trigonometric warping transformation is defined by Equation 1:
@@ 3.1.2 Cylindrical distortion transformation
Cylindrical warping deforms image content to simulate cylindrical surface distortions. This method mimics deformations caused by bounding line compression in books or reports, photographs of cylindrical surfaces, or curved distortions from photographing a document held aloft. Such deformations are highly prevalent in real-world scenarios and focus on practical requirements for document upright orientation. Consequently, the warping transformation is applied exclusively along the document’s vertical orientation (y-axis), without distorting the horizontal axis (x-axis). The cylindrical warping transformation is defined in Equation 2:

### v3

@@ Abstract
Table structure recognition is a key task in document analysis.However, geometric deformations in deformed tables weaken the correlation between content and structural information, thereby hindering downstream tasks’ ability to extract accurate content. To address this challenge, we propose the OG-HFYOLO model for fine-grained cell coordinate localization. The model integrates a Gradient-Orientation-Aware Extractor to enhance edge detection and introduces Heterogeneous Kernel Cross Fusion module to boost multi-scale feature learning, improving feature expression accuracy. Combined with a Scale-aware Loss function for better scale feature adaptation during training and mask-driven non-maximal suppression replacing traditional bounding-box suppression post-processing, the model achieves refined feature representation and superior localization performance. We further propose a data generator to address dataset limitations for fine-grained deformation table cell localization and construct the large-scale Deformation Wired Table (DWTAL) dataset. Experiments demonstrate that OG-HFYOLO achieves superior segmentation accuracy compared to all mainstream instance segmentation models on the DWTAL dataset. The dataset and the source code are open source: https://github.com/justliulong/OGHFYOLO.
@@ 1 Introduction
In the digital era of expanding information, tables serve as primary carriers of structured data, commonly conveying critical information in financial reports, educational materials, and scientific experimental results. Concurrently, the widespread adoption of scanning, photography, and other technologies has increased the complexity of table electronic document scenarios. Particularly under varying illumination, angles, and environmental conditions, scanned or photographed table images may exhibit diverse deformations. The physical deformations, such as bending, perspective distortion, and folds, pose significant challenges for table structure recognition technologies.
Table structure recognition technology, which reconstructs row-column topology and semantic information from table images, has advanced considerably through deep learning approaches [1]. While numerous end-to-end models exist and they can generate HTML or LaTeX table structure sequences directly from images [2, 3], their black-box nature makes the intermediate steps uncontrollable. Conversely, non-end-to-end strategies adopt a modular approach that decouples cell localization and structural inference, offering greater transparency in table structure recognition. However, one of the core task of such methods lies in precisely determining the spatial coordinates of table cells, which requires identifying the pixel-level positions of each cell within the image.
Current approaches to cell spatial coordinate localization can be broadly categorized into two main methods: One approach involves contour-based object detection, which can effectively identify cells with minor deformations. However, this method may omit critical cell content information with severe deformations, weakening the correlation between content and structure and complicating downstream tasks like content extraction. The other approach involves text box segmentation [4, 5], primarily used in wireless non-deformed tables but struggles with wired tables exhibiting geometric deformations. While keypoint detection (Figure 1(b)) partially addresses these challenges, it cannot fully resolve deformation-induced misalignment, leaving a critical gap in accurate coordinate acquisition for severely deformed wired cells.
To bridge this gap, we propose an instance segmentation-based framework (Figure 1(c)), which achieves pixel-level spatial localization and provides comprehensive structural information to support downstream tasks.However, two major challenges arise when applying instance segmentation to spatial localization of cells in deformed tables: First, the dense arrangement of table cells and shared contour lines between adjacent cells complicate the extraction of boundary information. Second, merged cells result in significant variations in cell scales, such as overly wide or narrow cells, further complicating cell spatial localization. Instance segmentation aims to achieve pixel-level object separation and semantic parsing, reflecting the exploration of balancing accuracy and efficiency in computer vision. Early research primarily utilized two-stage frameworks. A representative example is Mask R-CNN [6], which achieved highly accurate mask predictions in complex scenes through region proposal generation and feature extraction. However, due to the inherent latency of multi-stage computational processes, such methods struggle to meet real-time application demands. With the emergence of single-stage methods, researchers abandoned the region proposal mechanism, instead directly predicting object locations and masks via intensive feature map analysis. Nevertheless, single-stage models still encounter significant challenges in handling complex object outlines and differentiating dense objects, requiring feature expression mechanism innovations to enhance accuracy.
@@
Figure 1: Effect comparison of different methods: (1(a)) shows a deformed table under natural scene, (1(b)) demonstrates the detection results by contour corner-points, and (1(c)) presents the instance segmentation-based contour. As shown, instance segmentation maintains complete content in deformed cells, providing richer information for downstream tasks like content restoration.
@@
Beyond the challenges of the dense arrangement of objects and large-scale variations, current datasets for deformed table recognition tasks exhibit certain limitations. As per existing literature, while there are a few large-scale open-source datasets such as WTW[7] and TAL-OCR[8] that focus on deformed tables, their annotation granularity is insufficient to support the training of pixel-level segmentation models. To address this issue, we developed a data generator capable of creating new datasets from existing ones. Specifically, we selected mildly deformed tables from TAL-OCR and WTW, refined their labels into segmentation annotations, and used this annotated data with the data generator to produce a substantial amount of deformed table data, effectively supporting the training of instance segmentation models. Leveraging data from the open-source datasets TAL-OCR and WTW, along with 150 offline-collected and labeled images, we utilized the data generator to create a large-scale dataset. This dataset was split into two independent subsets based on their sources: DWTAL-s, comprising 8,765 simpler tables primarily from TAL-OCR, and DWTAL-l, consisting of 19,520 complex tables mainly expanded from WTW.
@@
To address the challenges of dense objectives and scale variations in the derived dataset, we first introduced the Gradient Orientation-aware Extractor (GOE). This module incorporates image texture gradient direction and strength into the YOLO-based network as learning features, enabling more comprehensive capture of cell boundary information. Additionally, during the feature fusion stage, the model employs the Heterogeneous Kernel Cross Fusion(HKCF), which integrates the bottleneck structure [9] and Heterogeneous Kernel Select Protocol (HKSP) [10]. This enhances the fusion of features across scales, improving the model’s detection performance under significant scale variations. Furthermore, we designed a Scale-aware loss function for segmentation tasks. This function adaptively assigns weights to various scale objects, enhancing the model’s robustness to scale variations. In the post-processing stage, traditional instance segmentation models typically employ bounding box-based non-maximal suppression (NMS). However, due to the dense arrangement and complex shapes of deformed cells, bounding boxes exhibit significant overlap. This causes boundary box-based NMS to incorrectly suppress correct objects. To address this issue, we introduced mask-based NMS, which directly performs suppression judgments using the intersection over union (IoU) of predicted masks. The main contributions of this work are summarized as follows:
@@
  • To address the issue of object omission in densely arranged cells of deformed tables, we propose the Gradient Orientation-aware Extractor (GOE), which enhances the boundary discrimination of densely arranged cells. The GOE integrates two key cues: gradient intensity and orientation. During feature extraction, directional filters decompose the image gradient field, quantifying intensity distribution in edge regions and preventing the isotropic diffusion of conventional convolution kernels along edges. Additionally, an orientation-aware attention mechanism is incorporated. This mechanism encodes gradient directions as channel components, guiding the deep network layers to focus on discriminative regions at deformation boundaries. We also propose the Heterogeneous Kernel Cross Fusion (HKCF) module for the feature fusion stage. This module dynamically adapts feature representation for multi-scale objects via the Heterogeneous Kernel Select Protocol (HKSP), deploying asymmetric convolution kernels in parallel. By integrating cross-layer feature interaction mechanisms, the HKCF explicitly captures distinct spatial patterns of horizontally wide and vertically narrow cells.
@@
  • We further design a scale-aware loss function that adaptively assigns weights to objects based on their scales. This enhances the model’s sensitivity to smaller deformed cells while reducing feature redundancy in larger cells. Furthermore, to prevent valid cells from being incorrectly suppressed during post-processing due to the complex shapes and dense arrangements of deformed cells, we replace conventional bounding box-based non-maximum suppression (NMS) with a mask-driven NMS post-processing operation.
@@
  • To address the scarcity of fine-grained annotated datasets for deformed tables, we designed a data generator that expands existing datasets. We first selected mildly deformed tables from existing datasets and converted their labels into finer-grained segmentation annotations. These curated samples were then processed by the generator to produce two datasets with varying difficulty levels and scales: DWTAL-s and DWTAL-l.
@@ 2.1 Spatial coordinate localization of table cells
Table cell spatial coordinate localization, a critical upstream task in table structure recognition, is essential for subsequent table parsing and structure recognition. Researchers have proposed various methods to address cell localization. For example, Prasad et al. [4] presented CascadeTabNet. This model treats table text boxes as objects and formulates text box detection as an instance segmentation task, extracting text box masks using Cascade Mask R-CNN. This marked the first application of instance segmentation algorithms in table structure recognition. However, CascadeTabNet’s experiments were restricted to non-deformable tables, and the impact of empty cells on downstream tasks was not extensively explored. Following a similar concept, Qiao Liang et al. [5] developed the LGPMA model. This model integrates global and local information via a soft feature pyramid and uses instance segmentation for text box detection. Unlike CascadeTabNet, LGPMA includes a specialized search algorithm for empty cells.
To address spatial coordinate extraction in deformable tables, Cycle-CenterNet [7] proposed a contour corner point-based object detection method that predicts four contour points by detecting cell centers. Although this method advances deformable table recognition, it has limitations. Severe boundary curvature increases the regression difficulty for corner points, and using coarse-grained object detection frameworks may result in the loss of critical cell information.
The challenges brought by deformed tables urgently require finer-grained techniques to preserve the key information of cells in the table. Driven by this requirement, our work integrates instance segmentation technology into spatial coordinate localization for deformable tables. By leveraging the pixel-level precision of instance segmentation, our approach achieves more refined spatial coordinate extraction, effectively addressing the challenges of cell localization in deformable table structures.
@@ 2.2 Existing dataset
The advancement of table structure recognition has been driven by numerous open-source datasets. Early datasets, such as UNLV [11] and ICDAR-2013 [12], were mainly designed to evaluate traditional methods. These datasets contain limited samples (typically fewer than 1,000 images) and lack spatial coordinate annotations for table cells. These limitations diminish their utility for modern deep learning-based modular table recognition pipelines. As deep learning requires increasingly larger datasets, researchers such as Xu Zhong and Minghao Li introduced large-scale datasets like PubTabNet [2] and TableBank [13]. However, these datasets focus on structural annotations via HTML or LaTeX sequences while overlooking spatial coordinate labeling. Similarly, large-scale datasets like FinTab [14] and SciTSR [15] include cell coordinates and row-column relationships but primarily consist of data derived from structured digital documents (e.g., PDFs or LaTeX-exported documents).
These datasets are highly standardized, making them unsuitable for deformable table recognition. Although ICDAR-2019 [16] attempted to address this gap by introducing scanned archival documents, its limited scale (3600 images) and deformation types are still inadequate. The CamCap [17] dataset, explicitly designed for rule-based algorithms, contains only 85 deformed tables, far too few to support data-driven deep learning models. Existing methods lacked robust solutions for tables in natural scenes with complex backgrounds and deformations until the WTW dataset [7] was introduced. This benchmark provides annotations specifically designed for such challenging scenarios. However, WTW uses coarse-grained spatial coordinate annotations based on four contour corner points, which are ineffective for severe deformations. Similarly, the education-focused TAL-OCR [8] dataset, collected from real-world photography, contains moderate deformations but lacks fine-grained instance segmentation-level spatial annotations for cells. None of these datasets adequately address the challenges of recognizing heavily deformed table structures.
Table 1: Benchmark Datasets for Table Structure Recognition Datasets
Dataset Year table Deformation
UNLV[11] 2012 558
ICDAR-2013[12] 2013 156
PubTabNet[2] 2020 568K
TableBank[13] 2019 417K
FinTab[14] 2020 89K
SciTSR[15] 2019 15K
ICDAR-2019[16] 2019 3.6K
CamCap[17] 2015 85 $\surd$
WTW[7] 2021 14K $\surd$
TAL-OCR[8] 2021 18K $\surd$
@@ 3 Data generator and derivative dataset
This section begins by outlining the implementation of the data generator, followed by a brief presentation of the DWTAL dataset, which is derived from this generator.
@@ 3.1 Data generator
Deep learning is inherently data-driven, with dataset scale and quality being critical factors in determining model performance. To generate a sufficiently large and high-quality dataset, we propose a data generator that creates distorted images and corresponding labels from existing images and annotations. The distortions are carefully designed to simulate real-world deformation scenarios. Specifically, the generator incorporates two distortion techniques: wave warping and cylindrical warping. Specifically, the generator incorporates two distortion techniques: wave warping and cylindrical warping. Additionally, it adjusts image brightness according to original illumination to mimic natural lighting variations in real-world photography. The subsequent three subsections detail the implementation of these two distortion techniques and the brightness adjustment strategy.
@@ 3.1.1 Wave distortion transformation
Wave warping is an image distortion technique based on trigonometric functions that creates wave-like deformations. This technique simulates distortions caused by attaching images to flexible curved surfaces (e.g., corrugated pipes or folded fabrics), which involve sine or cosine function-based warping, as well as human-induced distortions. Such warping introduces severe deformations, posing significant challenges for table structure recognition. The specific trigonometric warping transformation is defined as follows in Equation 1:
@@ 3.1.2 Cylindrical distortion transformation
Cylindrical warping distorts image content to simulate cylindrical surface deformations. This method mimics deformations caused by binding line compression in books or reports, photographs of cylindrical surfaces, or curved distortions from photographing a document held aloft. Such deformations are common in real-world scenarios and address the practical requirement for document upright orientation. Consequently, the warping transformation is applied exclusively along the document’s vertical orientation (y-axis) while preserving the horizontal axis (x-axis). The cylindrical warping transformation is defined as follows in Equation 2: